Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

2021-03-21

Observers:  Steve Williams, Timo Kravtsov , Claudia Gutiérrez, 
Support: Takashi Nagao
TAT: Tassilo Schweyer, Jesper Sollerman, Sheng Yang

EFOSC2 Calibrations: bias frames (EFOSC2), Flats (SOFI)

Weather:
Start: humidity 51%, wind 5 m/s, seeing 
Middle: humidity 26%, wind 7.9 m/s, seeing  1.24
End: humidity 24%, wind 9 m/s, seeing 1.38

Comments:
01:15: We start the night with a problem in the mirror (something with the detectors). Because of this, image analysis was not possible. The problem was not solved but we could start the observations.
05:49: The problems with the actorsA (in the mirror) continue. We changed to EFOSC and spent some time trying to solve the problem (doing some focus).

 UT Target 
 Type Setup  Slit Exposure (s) Airmass
Seeing 
Notes
00:00 SN2021dov Follow-up JH 205 1.2 1.15
00:10 AT2019hdl Follow-up JHK 3960 1.7 1.45  
01:25 SN2021egg Follow-up JH
2520 1.0 1.6
02:10 SN2021epv Follow-up JH 2520 1.0 1.26
03:00 SN2021dsb Follow-up JH 760 1.0 1.38  
03:15 SN2021fbd Follow-up JH 2520 1.1 1.38
03:58 SN2021fxy Follow-up JH
760 1.1 1.38
04:16 SN2021dlb Follow-up JH
2520 1.5 1.38
05:01 SN2021fqb Follow-up JH
2520 1.4 1.38
05:49 Changed to EFOSC2
Problems with the detectors.
06:35 AT2021gkh Classification g13 1.0
--


Aborted. Nothing visible in the acquisition image. The problems with the mirror continued. We did some Focus and jumped to the next object. The quality here was  bad.
06:50 SDSS1342 Follow-up g11+g16 1.0 1500 1.3 1.25
08:05 AT2021gln Classification g13 1.0 -- -- 1.38 Aborted. Nothing in the acquisition image.
08:15 AT2020sjw Follow-up g13 1.0 -- -- 1.38 Aborted. Object no longer visible.
08:32 LTT 7379 Standard g11, g13, g16 1.0/1.5 6*60 1.38
08:54 AT2021fzp Classification g13 1.0 900 1.38 Blue continuum
09:16 AT2021fpm Classification g13 1.0 300 1.2 1.38 SN Ia (>40 days)
09:34 AT2021fxf Classification g13 1.0 600
1.38SN Ic
Comments